OpenStudy (anonymous):

Factor completely. -2k - k³ - 3k² a) k(-k + 1)(k - 2) b) -k(k - 1)(k - 2) c) -k(k + 1)(k + 2)

OpenStudy (apples):

\[-2k-k^3-3k^2 = -k \cdot (k^2 + 3k + 2) = -k \cdot (k + 1) \cdot (k + 2) \]